OpenStudy (anonymous):

6x+3=y, How would i solve for "Y"

OpenStudy (anonymous):

it is solved for y already

OpenStudy (anonymous):

Is that the exact question you are looking at?

OpenStudy (anonymous):

Then how would i graph it?

OpenStudy (anonymous):

3 is the y-intercept, so (0,3) is a point on the graph. You just need to find one more point on the graph, plot it, and draw a line between the two points

OpenStudy (anonymous):

How would i do that?

OpenStudy (anonymous):

so if x = 1, then y = 6(1)+3 = 9, meaning (1, 9) is another point on the graph

OpenStudy (anonymous):

wait lemme check it really qucik

OpenStudy (anonymous):

Its wrong

OpenStudy (anonymous):

How would you solve for X

OpenStudy (anonymous):

what does your answer say? that isn't wrong, but i'm not sure what your teacher wants

OpenStudy (anonymous):

I described how to draw a picture (graph) of the line

OpenStudy (anonymous):

Okay this is how the question is written: "Graph each linear equation?" Then its gives me the problem which is : y=6x+3

OpenStudy (anonymous):

okay. you can use the method of plotting 2 points that I described to do that

OpenStudy (anonymous):

or you can plot the y-intercept and use slope

OpenStudy (anonymous):

Can you give me an example using this problem and using the y- intercepr and use slope method

OpenStudy (anonymous):

The slope is 6 (the number in from of the x term). This means that for every 1 unit we move in the positive x direction, the line goes up 6 in the y direction

OpenStudy (anonymous):

the y-intercept is (0, 3) so using slope if we move to x = 1, we need to add 6 to y. So another point is (1, 9)

OpenStudy (anonymous):

Then another would be (2, 15), etc..

OpenStudy (anonymous):

Wait lemme check again

OpenStudy (anonymous):

you just need to plot 2 points that are on the line to draw the line

OpenStudy (anonymous):

But in the back of my book the graph doesnt go that high, I only goes upto 5 and thats it.

OpenStudy (anonymous):

oh, ok

OpenStudy (anonymous):

I am still confused.

OpenStudy (anonymous):

The previouis problem was x-2y=6 and this one was the one after

OpenStudy (anonymous):

You can still plot two points and draw a line through them, the book just didn't use nice whole numbers to graph its line

OpenStudy (anonymous):

you could let y = 0 and find the x-intercept too

OpenStudy (anonymous):

0 =6x+3 x = -1/2 so (-1/2, 0) is another point on the graph

OpenStudy (anonymous):

you got Y already.